# Self Rewarding Self Improving

### v1

@@ Abstract
We demonstrate that large language models can effectively self-improve through self-judging without requiring reference solutions, leveraging the inherent asymmetry between generating and verifying solutions. Our experiments on Countdown puzzles and MIT Integration Bee problems show that models can provide reliable reward signals without ground truth answers, enabling reinforcement learning in domains previously not possible. By implementing self-judging, we achieve significant performance gains maintaining alignment with formal verification. When combined with synthetic question generation, we establish a complete self-improvement loop where models generate practice problems, solve them, and evaluate their own performance—achieving an 8% improvement with Qwen 2.5 7B over baseline and surpassing GPT-4o performance on integration tasks. Our findings demonstrate that LLM judges can provide effective reward signals for training models, unlocking many reinforcement learning environments previously limited by the difficulty of creating programmatic rewards. This suggests a potential paradigm shift toward AI systems that continuously improve through self-directed learning rather than human-guided training, potentially accelerating progress in domains with scarce training data or complex evaluation requirements.
@@ 1 Introduction
Reinforcement Learning (RL) has demonstrated remarkable potential for improving LLM performance. However, two significant challenges limit its broader application: the labor-intensive creation of training environments and the lack of sufficient data across many domains.
In this paper, we explore a novel approach: enabling models to self-improve through self-judging by leveraging the generator-verifier gap. We investigate whether models can provide reliable reward signals without access to ground truth answers—simply by assessing the correctness of solutions directly. This approach could dramatically expand the scope of possible environments for RL training.
A key innovation of our work is leveraging the generator-verifier gap—the principle that verifying correctness is often simpler than generating solutions. We explore whether this inherent asymmetry enables models to provide reliable reward signals without access to ground truth answers.
Self-judging provides an extremely promising direction for expanding reinforcement learning to previously inaccessible domains. Numerous sophisticated tasks—such as engineering precise mechanical components like M4 bolts—present significant obstacles when developing reinforcement learning environments, as creating thorough design validation systems requires substantial effort and resources. Similarly, advancing AI agents that can effectively navigate computer interfaces has been hindered by the complexity of constructing suitable training environments. Self-judging could potentially solve this bottleneck by allowing models to evaluate performance in domains where defining programmatic reward functions are difficult.
This approach is particularly valuable when paired with synthetic question generation. A model creating its own practice problems and then evaluating its solutions establishes a complete self-improvement loop, dramatically reducing the need for extensive human-annotated datasets. This creates opportunities for autonomous skill development in specialized domains where labeled examples are scarce but verification criteria exist.
We demonstrate our approach through three experimental setups. First, using Qwen 2.5 7B, we explore training dynamics on the "Countdown" arithmetic task, where models must construct expressions to reach target values. While susceptible to reward hacking, we show that with appropriate prompting, self-judging models can match formal verification performance. The robustness of different judging prompts varies significantly, with some configurations maintaining alignment with formal verification throughout training while others diverge as agents discover exploits.
Building on our LADDER framework to generate synthetic training data, we demonstrate that LLMs functioning as judges provide stable reward signals for integration tasks, enabling a 7B model to surpass GPT-4o performance without ground truth access. By implementing a system where models generate their own practice problems, solve them, and then evaluate the correctness of their solutions, we establish a complete self-improvement loop without requiring external verification. This proves particularly effective for mathematical reasoning tasks, where we achieved an 8% improvement with Qwen 2.5 7B over baseline. We further investigate weak-to-strong generalization by using GPT-4o as a judge with Deepseek Distilled Qwen 2.5 7B as the agent, achieving performance approaching O1 levels on challenging integration problems.
Our approach reveals that models can generate synthetic tasks and reliably evaluate their own performance in certain domains, pointing toward a new paradigm of autonomous self-improvement. This has profound implications for AI development, potentially enabling systems that continuously identify weaknesses, generate appropriate practice, and improve through self-directed learning cycles. Such autonomy could dramatically accelerate progress in domains where traditional supervised learning is limited by data availability or annotation costs.
2 Related Work
Our approach builds upon several research streams in self-improvement and LLM training. Meta’s Self-Rewarding Language Models work demonstrated that models can generate preference data using an LLM as a judge with Direct Preference Optimization [Yuan et al., 2025] that led to significant performance improvements on benchmark tasks. Unlike Meta’s preference-based approach, our method focuses specifically on mathematical reasoning domains with verifiable solutions.
We also extend our previous work [Simonds and Yoshiyama, 2025] LADDER framework, which enabled improvements on mathematical benchmarks through recursive problem decomposition and demonstrated capabilities for building self-improving systems for reasoning tasks. Our work specifically builds on LADDER’s synthetic question generation capabilities, but implements LLM as a judge rather than a numeric verifier.
Our research connects to weak-to-strong supervision literature, where [Burns et al.,] demonstrated that larger models supervised by smaller models can achieve strong performance on tasks the smaller model couldn’t solve consistently. This relates to concepts from iterated amplification [Christiano et al., 2018] and addresses alignment concerns raised[Hubinger et al., 2021], that warn about reward hacking in learned optimizers. Our approach mitigates these risks by using verifiable signals rather than opaque weak supervision.
We also build upon code generation frameworks like CodeRL [Le et al., 2022] that use unit test outcomes as reward criteria and verification-based learning from solved examples. Unlike approaches that rely solely on external verification systems, our method establishes a framework where models practice on self-created exercises with reliable self-judgment, enabling continuous improvement in domains where defining programmatic reward functions is challenging.
@@ 3 Methodology
3.1 Self Judging Setup
Our self-judging approach employs an LLM as a reward model without access to ground truth answers. The judge LLM receives the agent’s solution and the initial question, and produces a binary "correct" or "incorrect" evaluation. We implemented this as an offline system where the judge model remained fixed throughout training to ensure stable reward signals. To minimize reward hacking, we constrained the judge’s input to only the agent’s solution within designated answer tags, preventing manipulation through extraneous explanations or commentary that might bias evaluation.
For each experimental domain, we designed specific prompting strategies based on preliminary analysis of judge reliability. For RL training, we implemented Group Relative Policy Optimization (GRPO). Training was conducted with a batch size of 64 across all experiments.
3.2 Experiments
We conducted experiments across two primary tasks to evaluate the effectiveness of LLM self-judging in reinforcement learning.
The first task, Countdown, presents a mathematical puzzle where the model must combine a set of given numbers using basic arithmetic operations to reach a target number. For example, given numbers 2, 3, 5, 7 and target 19, a valid solution might be (7 × 3) - 2. This task requires both computational accuracy and strategic thinking to explore the solution space efficiently.
The second task utilizes questions from the 2025 MIT Integration Bee qualifying examination—a prestigious competition attracting undergraduate and graduate students with strong mathematical backgrounds. The qualifying exam consists of 20 integration problems with a typical qualifying threshold of 73%, though most participants score between 15-30%, with 50-73% considered strong performance.
For our training set, we leveraged variants generated using our LADDER framework to address the limited size and sparse difficulty distribution of the original exam questions. The LADDER approach generates synthetic integration problems with controlled difficulty levels, providing a more robust training dataset while maintaining the mathematical characteristics of the original MIT problems. We generated 9000 variants for our training set
For evaluation, we employ a formal verification system. In the Countdown task, solutions are automatically checked for correctness by verifying that the arithmetic operations on the provided numbers yield the target value. For the MIT Integration problems, we implement a two-stage evaluation process: during training, models are required to output solutions in SymPy format, enabling programmatic verification of correctness. For final results, we conduct manual assessment to account for mathematically correct solutions that may contain formatting errors in the SymPy syntax. This manual review is especially important since our base model show poor initial accuracy in producing valid SymPy syntax, allowing us to differentiate between genuine improvements in integration problem-solving abilities versus mere enhancements in syntax formatting.
@@ 4.1 Countdown
We begin by examining Countdown, in which given a set of numbers and a target value, the task is to construct an arithmetic expression that evaluates to the target value. We choose this task to start with, since it possesses two relevant properties: First, previous work has empirically shown that training a LLM on this task using RL elicits reasoning abilities [Pan et al., 2025]. Second, verifying a potential solution is intuitively simpler than generating one from scratch.
An agent being a robust self-judge, should lead to similar training dynamics when replaced by a formal verifier implementing the same algorithm. However, designing a prompt for a LLM judge, that acts as a reliable proxy for the formal reward function its a challenge. Note that language can be ambiguous, and in addition there is no explicit control over the sampled tokens that implement the proxy reward function, therefore making reward signals non-deterministic and less reliable. These opens the possibility for an agent to find strategies that hack the proxy reward.
@@
Figure 2 illustrates the training dynamics of a Qwen 2.5 7B model across the four prompts. Early in training, all prompts led the True Negative Rate (TNR) of the self-rewarding agent approach its upper limit. However as training progressed, the policy gradient optimization algorithm found a strategy to reward hack the less explicit prompts, which resulted in a substantial drop in the TNR, except for the last prompt. Ultimately the agent’s mean self-reward diverged from the mean formal reward, leading to poor evaluation metrics. Examples of the hacks found by the agent for all three prompts, and the mean formal reward on the evaluation set can be seen in Appendix A.
Figure 2: Self-rewarding training dynamics of Qwen 2.5B 7B for the different prompts. Both the TNR and True Reward are computed using a formal verifier.
[…]
The previous experiments demonstrate that a single model can act both as a generator and as a verifier, enabling self-improvement through self-verification in a domain where there exists a generator-verifier gap.
@@
4.2 Simple Integration
Refer to caption Refer to caption
Figure 4: Qwen 2.5 7B Simple Integration Self Judging Dynamics
Before tackling the MIT Integration Bee, we first explored judge dynamics on a simpler integration dataset with known solutions. We specifically examined Qwen 2.5 7B self-judging its own solutions.
Our analysis revealed that despite relatively constant error rates (False Negative and False Positive rates of approximately 10%), the model consistently improved performance on both test and training sets (see Figure 4). The initial low validation score was primarily attributed to output formatting issues with Qwen models. When accounting for formatting issues actual accuracy was 54%. This highlights an advantage of LLM-based judges: they are less sensitive to formatting variations than formal verification tools. We note this may introduce some noise in the False Positive measurements due to models correctly marking solutions that have incorrect formatting and hence couldn’t be assessed using our formal verification tools.
Interestingly, we observed that our judge model initially output higher True Positive rewards but grew significantly slower than the validation score improvement. This divergence is particularly notable given that both test and training sets were sampled from the same base dataset with equivalent difficulty levels.
Analyzing the error rates over time, we found that the False Positive Rate (FPR) gradually increased, primarily due to decreasing True Negatives. In contrast, the False Negative Rate (FNR) remained relatively constant. This stability masked an underlying dynamic where both False Negatives and True Positives grew proportionally.
Perhaps the most significant finding is that despite these noisy judge signals, the model continued to improve its integration-solving capabilities. This suggests that even imperfect feedback mechanisms can drive meaningful performance gains in mathematical reasoning tasks.
Refer to caption Refer to caption
Figure 5: Qwen 2.5 Simple Integration Self Judging Dynamics
4.3 Self Judging
Figure 6: Qwen 2.5 7B + Qwen 2.5 7B Judge
We conducted an experiment exploring complete self-improvement through Qwen 2.5 7B both generating and judging integration problems. In this setup, the model generates synthetic integration questions via our LADDER framework and then evaluates its own solutions, creating a closed self-improvement loop without external validation or tool use.
Our results showed steady performance improvement, rising from an initially low score (hampered by formatting issues with SymPy output) to a performance of 43% on the MIT Integration Bee evaluation set. This demonstrated an 8% improvement over the base model surpassing GPT 4o (42%). While the model eventually plateaued, this represents substantial improvement through purely self-directed learning.
This experiment demonstrates a powerful paradigm: a single model architecture serving as both problem generator, solver, and judge. The ability to improve significantly without external feedback suggests models can identify their own weaknesses, practice appropriately, and verify their improvement—key components of autonomous learning systems. Notably, since we did not update the judge model (using the base Qwen 2.5 7B throughout), the performance plateau likely stemmed from judge limitations. Future work should explore online methods where both judge and generator models improve simultaneously, potentially overcoming such bottlenecks.
Qwen 7B GPT-4o Qwen 7B+RL Judge 20202020303030304040404050505050PerformanceQwen 2.5 7B with Qwen 2.5 7B Judge Qwen DS Distilled Qwen DS Distilled + GPT-4o Judge o14040404050505050606060607070707080808080PerformanceQwen 2.5 DS Distilled 7B with GPT 4o Judge
Figure 7: Performance comparison of baseline and RL-enhanced models
4.4 Weak to Strong Supervision
Figure 8: Qwen 2.5 7B + GPT 4o Judge training progress
We examined the effectiveness of using a stronger model (Qwen 2.5 7B Deepseek Distilled) as a generator with a weaker model (GPT 4o) as a judge. This setup tested whether a model with superior reasoning capabilities could provide effective reward signals despite having lower baseline performance on the specific task—GPT-4o scores 42% on the MIT Integration Bee benchmark compared to Qwen’s 50%.
Our results were striking: with GPT-4o as judge, the Qwen agent achieved average performance of 67%(Figure 7) on the validation set. This represents substantial improvement compared to the baseline and demonstrates that a judge model doesn’t necessarily need domain-specific expertise exceeding the agent to provide effective feedback. The model was still showing consistent improvement when we concluded our experimental runs, suggesting further gains might be possible with additional training. However, we made the practical decision to end experiments at this point due to API pricing constraints.
GPT-4o proved significantly less sensitive to prompt variations than Qwen models. The larger model showed greater resistance to reward hacking attempts, maintaining alignment between its judgments and mathematical correctness.
These findings highlight a promising dynamic: while small LLM judges require careful prompt engineering to prevent exploitation, stronger models naturally serve as more reliable judges with less engineering overhead—indicating potential for increasingly autonomous and stable self-improvement systems as model capabilities advance. However, it remains an open question regarding the scaling laws at play—whether models’ ability to exploit judges improves faster than judges’ ability to detect exploitation as capabilities increase. This competitive dynamic will likely be a crucial factor in determining the long-term viability and stability of self-supervised improvement systems.

### v2

# RLSR: Reinforcement Learning from Self Reward

@@ Abstract
Large language models can generate solutions to complex problems, but training them with reinforcement learning typically requires verifiable rewards that are expensive to create and not possible for all domains. We demonstrate that LLMs can effectively self-improve through self-judging without reference solutions, leveraging the inherent asymmetry between generating and verifying solutions. Our experiments show that models can provide reliable reward signals without ground truth answers, enabling reinforcement learning in domains where verifiable rewards are impractical. By implementing self-judging across Countdown puzzles and integration problems, we achieve performance comparable to formal verification without ground truth solutions. Most notably, Qwen 2.5 7B DeepSeek Distilled trained with self-rewards qualifies for the prestigious MIT Integration Bee competition, performance through self-supervised improvement. When combined with synthetic question generation, we establish a complete self-improvement loop where models generate practice problems, solve them, and evaluate their own performance without any external validation. Our findings demonstrate that LLM judges can provide effective reward signals for training, unlocking reinforcement learning in countless domains previously limited by reward engineering challenges. This work represents a significant step toward autonomous AI systems that continuously improve through self-directed learning rather than human-guided training, potentially accelerating progress across domains where training data is scarce or evaluation is complex.
@@ 1 Introduction
Reinforcement Learning (RL) has proven transformative for improving large language model capabilities, from mathematical reasoning to code generation. Yet its broader adoption faces two fundamental barriers: the engineering complexity of creating verifiable reward functions and the scarcity of training data across specialized domains. Current approaches require substantial human effort to design training environments, implement reward functions, and guard against reward hacking—limiting RL to well-resourced domains with clear programmatic verification. In this work, we explore a paradigm shift: can models reliably improve themselves through self-evaluation, without access to ground truth answers?
The key insight enabling our approach is the generator-verifier gap—the fundamental asymmetry where verifying a solution’s correctness is computationally simpler than generating it from scratch [1, 2]. This principle manifests across numerous domains: validating an M4 bolt design against specifications is straightforward compared to designing one, checking a mathematical proof is easier than discovering it, and evaluating whether an AI agent successfully completed a computer task is simpler than programming the agent. While this asymmetry is well-established theoretically, its practical application to self-improving systems remains largely unexplored. By having models judge their own outputs, we can potentially transform any domain with this asymmetry into a viable RL training environment.
Our core contribution is demonstrating that LLMs can serve as reliable judges of their own performance, enabling self-improvement without external verification. We show that carefully designed self-judging systems can match the training dynamics of formal reward functions while dramatically expanding the scope of trainable domains. Crucially, when combined with synthetic question generation, this creates a complete autonomous learning loop: models generate their own practice problems, attempt solutions, and evaluate their performance—all without human intervention or ground truth data.
We validate this approach through systematic experiments across three domains of increasing complexity. In Countdown arithmetic puzzles, we identify critical challenges around reward hacking and develop robust prompting strategies that prevent exploitation while maintaining reliable training signals. Through ablation studies on simple integration problems, we quantify the generator-verifier gap across different models and explore how judge dynamics affect learning outcomes. Most compellingly, we then demonstrate that Qwen 2.5 7B[3] trained entirely through self-judging achieves performance qualifying for the MIT Integration Bee—a prestigious mathematics competition—surpassing through purely self-supervised improvement.
The implications extend far beyond our specific experiments. Self-judging fundamentally changes the economics of applying RL to new domains by eliminating the need for custom training environments and programmatic rewards. Domains previously considered impractical for RL—from engineering design to creative writing evaluation—become accessible. More profoundly, this work provides evidence that AI systems can engage in genuine self-improvement, identifying their weaknesses, generating targeted practice, and iterating without human guidance. While challenges remain, particularly around preventing reward exploitation and ensuring robust evaluation, our results suggest a path toward truly autonomous learning systems.
Key Contributions
  • We demonstrate that large language models (LLMs) can self-judge their own outputs without ground-truth labels, achieving training dynamics comparable to formal verification systems.
  • We identify—and resolve—critical challenges in self-judging, especially reward hacking in smaller models, via systematic prompt engineering.
  • We establish a complete self-improvement loop that combines synthetic question generation with self-evaluation, enabling fully autonomous skill development.
  • We apply our techniques to qualify for MIT Integration Bee problems through pure self-supervision, using Qwen 2.5 7B DeepSeek Distilled[4].
@@ 2 Methodology
2.1 Self-Judging Framework
Our approach centers on using LLMs as reward models for their own outputs, without access to ground truth answers. For each generated solution, we prompt the model to provide a binary evaluation—“correct” or “incorrect”—based solely on examining the solution attempt and the original problem. This creates a self-contained feedback loop where the same model architecture serves as both policy (generating solutions) and reward function (evaluating them). For all experiments unless otherwise stated we use an offline judge that doesn’t update during training.
For RL training, we employ Group Relative Policy Optimization (GRPO)[5] with a batch size of 64 and rollout of 8 for all experiments. We use a learning rate of $1\times 10^{-6}$, KL loss coefficient of 0.001, and maximum response length of 2048 tokens (extended to 8096 for MIT Integration experiments to accommodate complex solutions). We provide an additional 0.1 format reward for solutions formatted correctly. For our integration experiments this includes outputting valid SymPy inside the answer tags
2.2 Experimental Domains
We selected three domains that systematically explore different aspects of self-judging:
Countdown tests whether self-judging is feasible at all. This arithmetic puzzle requires combining given numbers to reach a target—simple enough to analyze failure modes, yet complex enough to benefit from RL-induced reasoning improvements. We synthetically generated a dataset of 4-digits countdown problems (sampled between 0 and 100) that must be combined using basic arithmetic operations to reach a target value. Our dataset consists of 2,000 training problems and 200 test problems. Prior work has shown RL enhances reasoning on this task, making it an ideal testbed. Here we focus on identifying and preventing reward hacking through iterative prompt design.
Simple Integration allows us to quantify the generator-verifier gap—the asymmetry between the difficulty of solving problems versus checking solutions. We constructed a dataset of college-level integration problems synthetically generated by GPT-4o, consisting of 2,000 training problems and 200 test problems. By requiring solutions in a structured format (answer tags with valid SymPy expressions), we can compare self-judging against formal verification. This domain lets us systematically study how judge quality affects downstream performance improvements while maintaining consistent difficulty and style across the dataset.
MIT Integration Bee demonstrates real-world effectiveness. The MIT Integration Bee is an annual mathematics competition that attracts top undergraduate and graduate students. The qualifying examination is notoriously challenging—in typical years, only 16 students achieve the 73% qualifying threshold out of all participants. We evaluate on the 2025 competition problems, representing genuinely difficult integration challenges that test the limits of mathematical reasoning. Each years exam only has 20 questions. To create sufficient training data, we augment the limited competition dataset with 9,000 synthetic variants generated via our LADDER framework using Qwen 2.5 7B DeepSeek distilled, maintaining problem characteristics while providing adequate coverage for RL training. We filter out problems from our training set that the base model solves pass@1
2.3 Judge Prompting and Evaluation
For integration problems, we employ a carefully designed prompt that leverages the generator-verifier gap:
Please check if the following is a valid function: {}. If it is, differentiate it and determine if it is functionally equal to {}. Output <JUDGE_SCORE>1</JUDGE_SCORE> if they are equal. Output <JUDGE_SCORE>0</JUDGE_SCORE> if they are not equal or if it is not a valid function. Ignore constants of integration.
This prompt instructs the judge to first verify syntactic validity, then apply symbolic differentiation—a significantly easier task than integration—to check correctness. We selected this prompt based on preliminary experiments showing consistent performance across different model scales.
2.4 Evaluation Framework
Each domain employs formal verification to assess true performance, enabling us to measure how well self-judging approximates ground truth:
  • Countdown: Solutions are programmatically verified by executing the arithmetic operations and checking if they yield the target value
  • Integration tasks: We verify solutions using symbolic differentiation—if the derivative of the proposed answer matches the original integrand, the solution is correct
A consistent challenge in the integration tasks was capturing model outputs in a parseable format. To address this, we required models to output their answers in SymPy syntax, which provided more consistent formatting than raw mathematical notation. However, this introduces a confounding factor: models that fail to produce valid SymPy syntax are marked incorrect, even if their mathematical reasoning is sound. While we attempted to be generous by also accepting some LaTeX formatting, parsing failures remained a persistent issue.
This format requirement means that initial performance improvements often reflect the model learning proper syntax rather than enhanced mathematical ability. We typically observed rapid early gains as models learned to output valid SymPy expressions, followed by slower improvements that more likely represent genuine advances in integration skill. This distinction is important when interpreting our results—early training dynamics may overstate actual mathematical reasoning improvements.
@@ 3.1 Countdown
We begin by examining Countdown, in which given a set of numbers and a target value, the task is to construct an arithmetic expression that evaluates to the target value. We choose this task to start with, since it possesses two relevant properties: First, previous work has empirically shown that training an LLM on this task using RL elicits reasoning abilities [6]. Second, verifying a potential solution is simpler than generating one from scratch. All experiments were run using Qwen 2.5 7B Instruct as both the agent and judge.
To establish the generator-verifier gap for this domain, we first measured baseline performance on our synthetic dataset of 2,000 training problems. Qwen 2.5 7B Instruct achieved approximately 36% accuracy on generating correct solutions. To evaluate the model’s judging capabilities, we constructed a balanced verification dataset using GPT-4o, comprising equal proportions of correct and incorrect solution attempts. When tasked with classifying these solutions, the same model achieved 96% accuracy—demonstrating a substantial 62 percentage point gap that validates the feasibility of self-judging for this task.
For a robust self-judge, the agent should lead to similar training dynamics when replaced by a formal verifier implementing the same algorithm. However, designing a prompt for a LLM judge, that acts as a reliable proxy for the formal reward function is a challenge. Note that language can be ambiguous, and in addition there is no explicit control over the sampled tokens that implement the proxy reward function, therefore making reward signals non-deterministic and less reliable. These opens the possibility for an agent to find strategies that hack the proxy reward.
@@
Figure 2 illustrates the training dynamics of a Qwen 2.5 7B model across the four prompts. Early in training, all prompts led the True Negative Rate (TNR) of the self-rewarding agent approach its upper limit. However as training progressed, the policy gradient optimization algorithm found a strategy to reward hack the less explicit prompts, which resulted in a substantial drop in the TNR, except for the last prompt. Ultimately the agent’s mean self-reward diverged from the mean formal reward, leading to poor evaluation metrics.
Figure 2: Self-rewarding training dynamics of Qwen 2.5 7B for the different prompts. Both the TNR and True Reward are computed using a formal verifier.
A comparison of the training dynamics of a self-judging agent using the most robust prompt, and an agent using a formal reward function is displayed in Figure 3. A consistent performance gain of approximately $20\%$ over the baseline was observed in both configurations. The mean reward curves also exhibited analogous trends across these conditions. Furthermore, the mean response length increased at similar rates in both scenarios, a phenomenon attributed to the emergence of search abilities within the agent chain of thought.
@@
3.2 Simple Integration Ablations
To understand when and why self-judging succeeds, we systematically explore the generator-verifier gap—the fundamental asymmetry that makes verification easier than generation. We examine three judge configurations (Llama 3.2 3B, Qwen 2.5 7B, and DeepSeek R1) and compare their downstream training performance against a formal verification baseline.
3.2.1 Quantifying the Generator-Verifier Gap
We first establish empirical evidence for the generator-verifier gap in integration tasks. Our hypothesis is straightforward: if models find verification substantially easier than generation, they should be more accurate at checking solutions (via differentiation) than producing them (via integration).
To test this, we evaluated each model on two complementary tasks using our simple integration dataset:
  • Generation (Integration): Given an integrand, produce its antiderivative
  • Verification (Differentiation): Take the output antiderivative (generated using symbolic system) and apply derivative to generate the original integrand
For both tasks, we employed robust parsing augmented with manual evaluation to handle formatting variations while maintaining evaluation integrity. This dual approach ensures we capture true mathematical capability rather than just syntax compliance. We test models on the test set of our simple integration dataset
Figure 4 presents our results. Across all three models, we observe a consistent pattern: verification accuracy substantially exceeds generation accuracy. Llama 3.2 3B achieves 47% accuracy on differentiation despite 1% success on integration—demonstrating that even models incapable of solving problems can potentially judge solutions. Qwen 2.5 7B shows a similar pattern (88% verification vs 46% generation), while DeepSeek R1 nearly saturates both tasks but maintains a small gap. The existence of this gap suggests that even models struggling with integration can potentially provide meaningful reward signals by leveraging their superior verification abilities.
Llama 3.2 3BQwen 2.5 7BDeepSeek R100.20.20.40.40.60.60.80.811ModelAccuracyGenerator-Verifier Gap Across ModelsIntegration (Generation)Differentiation (Verification)
Figure 4: Generator-verifier gap demonstrated through integration vs differentiation accuracy. All models show substantially higher performance on verification (differentiation) compared to generation (integration), with gaps ranging from 47 percentage points (Llama 3.2 3B) to 1 percentage point (DeepSeek R1). Error bars represent standard error.
3.2.2 Judge Quality and Downstream Performance
Having established that models possess asymmetric verification capabilities, we now examine how judge quality translates to actual training performance. We trained Qwen 2.5 7B Instruct using four different reward sources: three self-judge configurations (Llama 3.2 3B, Qwen 2.5 7B, and DeepSeek R1) and symbolic verification as our gold standard baseline.
This experiment directly tests our central hypothesis: can models with sufficient verification accuracy enable effective self-improvement, despite being unable to solve many problems directly? By comparing judges with varying generator-verifier gaps (from 47% to near-perfect verification accuracy), we can identify the minimum judge capability required for meaningful learning.
Figure 5(a) presents the training curves, revealing a clear relationship between judge verification accuracy and downstream performance. When using Llama 3.2 3B as judge, with its 47% verification accuracy, the agent plateaus at approximately 48% test accuracy—barely above the judge’s own verification performance. This suggests the model primarily learns output formatting without genuine mathematical improvement. The judge’s limited verification ability creates a performance ceiling that prevents the agent from surpassing the judge’s own capabilities.
In contrast, Qwen 2.5 7B’s 88% verification accuracy enables substantial improvement, with the agent reaching 75% accuracy after 150 training steps. This demonstrates that judges need not be perfect—88% verification accuracy proves sufficient for meaningful self-improvement, despite the judge solving only 46% of problems directly. The dramatic performance difference between these two self-judges illustrates that judge quality is the critical factor determining training success.
At the upper end, DeepSeek R1 achieves performance comparable to symbolic verification accuracy. This convergence suggests that beyond a certain threshold, further improvements in judge accuracy yield diminishing returns—a near-perfect self-judge with 99% verification accuracy can fully substitute for formal verification. The symbolic verification baseline establishes this 92% as the upper bound, likely limited by the agent’s inherent capabilities rather than reward signal quality. We stop R1 runs early due to latency of API.
These results validate our generator-verifier framework: effective self-judging requires sufficient verification accuracy but not perfect generation ability. When judges are strong enough, self-improvement can approach the theoretical maximum achieved with perfect reward signals, opening the door for self-supervised learning in domains where formal verification is impractical or impossible.
(a) Qwen 2.5 7B Instruct agent with different judges (test-set accuracy)
(b) Qwen 2.5 7B Instruct – Updated vs Base Judge (test-set accuracy)
3.2.3 Judge Updates
The previous experiments assumed a fixed judge throughout training which remains at its initial capability while the agent potentially surpasses it. We hypothesized that periodically updating the judge to match the agent’s improved weights could enable continued learning beyond the initial judge’s limitations allowing us to overcome plateaus in agent performance.
To test this, we compared two training configurations using Qwen 2.5 7B as both agent and judge. In the baseline condition, the judge remained fixed at its initial weights throughout training. In the experimental condition, we updated the judge weights to match the agent’s current weights after 50 training steps, creating a self-evaluation system that evolves with the agent’s capabilities.
Figure 5(b) shows the training curves for both configurations. The fixed judge setup plateaus at approximately 70% accuracy, consistent with our earlier findings that agents struggle to substantially exceed their judge’s capabilities. In contrast, the updated judge configuration continues improving after the weight update, ultimately reaching 75% accuracy—a 7% relative improvement over the fixed baseline.
To understand why updating enabled continued improvement, we evaluated the verification accuracy of the updated judge model. When tested on our benchmark, the judge after 50 steps of training achieved approximately 93% accuracy on taking the derivative—a 5 percentage point improvement from the initial 88%. This enhanced verification capability explains the continued learning: the agent had developed better judgment abilities during training, which when leveraged for self-evaluation, enabled it to push beyond the original judge’s limitations.
This result reveals an important dynamic in self-judging systems: training can simultaneously improve both generation and verification capabilities. The 5% improvement in verification accuracy, while modest, proved sufficient to unlock additional performance gains. This suggests that even small improvements in judge quality can have meaningful impacts on downstream training outcomes, particularly when the judge already operates at high accuracy levels.
However, this approach introduces new considerations. Updating judges during training potentially reduces evaluation stability and could allow agents to exploit evolving reward signals. We tested a configuration where the judge updates after every training step, making the agent and judge identical throughout training. This led to rapid training collapse—the model’s performance degraded as it learned to exploit its own evaluation biases, creating a feedback loop where increasingly poor solutions received high rewards.
This failure mode highlights the delicate balance required in self-judging systems. While our single update after 50 steps proved beneficial, continuous self-evaluation appears fundamentally unstable. The model needs some distance between its generation and verification roles to maintain honest evaluation. Even with our successful single-update approach, performance eventually plateaus, suggesting inherent limits to self-improvement through self-evaluation. Future work should explore intermediate update frequencies to find the optimal balance between leveraging improved verification capabilities and maintaining evaluation integrity.
3.3 MIT Integration Bee: Real-World Effectiveness
0202040406060808010010012012014014016016002020404060608080100100Training StepsAccuracy (%)MIT Integration PerformanceQwen 2.5 7B DS + Self-JudgeQualification Score
Figure 6: Test set accuracy progression through training run
To demonstrate the practical impact of self-judging, we evaluated whether our approach could enable a model to qualify for the MIT Integration Bee—an annual mathematics competition where only 16 out of hundreds of participants achieve the 73% qualifying threshold. This represents a genuinely challenging benchmark that tests the limits of mathematical reasoning.
We trained Qwen 2.5 7B DeepSeek Distilled using self-judging with judge updates every 50 steps. The model evaluated its own integration attempts using our differentiation-based verification prompt, creating a complete self-improvement loop without access to ground truth solutions.
Figure 6 shows the training progression. Initial performance was poor due to formatting issues—the model struggled to output valid SymPy expressions required by our evaluation framework. Once these formatting challenges were resolved around step 50, accuracy improved from 13% to 51%. Performance then continued to climb steadily, crossing the 73% qualification threshold at approximately 160 steps and reaching 75.5% by step 170.
These results demonstrate that self-judging can achieve human-competitive performance on prestigious mathematics competitions through pure self-supervision. The model qualified for the MIT Integration Bee without any ground truth solutions or external verification—relying entirely on its ability to check its own work via differentiation. This validates that self-judging is not merely a theoretical curiosity but a practical training method capable of producing state-of-the-art results on challenging real-world tasks.